Philipp Maier

Street 15 1234AB G-Town The Netherlands

Employer Research London E-Mail: application@employer.com

Date

13th March 2001

Application for a position in the Research Department

Dear Mrs. Abc,

Hereby I wish to apply for a position at the Research Department of Employer Research in London.

Following a general interest in economic affairs, I studied General Economics in Germany. I finished my studies in three years, graduated as best of class and received the Albert-Einstein award for re-inventing the weel. Etcetcetc.

Working as an intern at the Beployer Research company I realized that I am truely interested in economic research. I have decided to apply at Employer Research in London, as …

Please do not hesitate to contact me if details regarding my CV or this application letter are unclear.

Yours sincerely,

Philipp Maier

Full CV attached

Phone: +31 50 123.4556
E–Mail: pmaier@yahoo.com
HTTP: www.philipp-maier.de